# First operation of LArTPC in the stratosphere as an engineering GRAMS balloon flight (eGRAMS)

### v1

@@ 1.2 Cosmic-Ray Antiparticles
Various cosmic particles, including antiparticles such as antiprotons and positrons, are incident on the Earth from space. The fluxes of antiprotons have been detected and measured by numerous experiments [5, 6, 7]. The majority of the measured antiproton flux is explained by secondary production in which primary cosmic rays such as protons colliding with interstellar gas: $p_{\mathrm{CR}}+p_{\mathrm{ISM}}\rightarrow\bar{p}+X$. However, a potential excess exists in the low energy region around 1 GeV. While still debated, this excess could be potentially described by the annihilation of dark matter (DM) with a mass around 50 GeV [8]. Despite the high-statistics antiproton measurements, cosmic-ray antideuterons have never been observed. Currently, the only upper limit on the antideuteron flux is placed by BESS Polar-II experiment [9]. When considering the dark matter model that describes the potential antiproton excess, the predicted antideuteron flux at energies below 0.2 GeV/n from dark matter annihilation is significantly larger (S/N $>$ 100) than the expected astrophysical secondary flux. This difference arises due to the kinematic limitations on the secondary production. In contrast, dark matter annihilation lacks these limitations, allowing the production of low-energy antideuterons. Consequently, the search for antideuterons has gained attention as a potential smoking gun for indirect DM search.
@@ 1.3 The GRAMS Experiment
The GRAMS (Gamma-Ray and AntiMatter Survey) experiment is a pioneering next-generation balloon/satellite experiment. It is the first project to utilize a LArTPC for simultaneous astrophysical observations of cosmic MeV gamma-rays and indirect dark matter searches using cosmic antiparticles [13]. Figure 1 illustrates the conceptual design of GRAMS and the detection principle of a LArTPC. As shown in Figure 1, the GRAMS detector consists of a LArTPC surrounded by two layers of Time-of-Flight (ToF) scintillators. For charged particles, the ToFs can provide velocity information of the incoming particle and for gamma rays, it can act as a veto. When a charged particle enters the LArTPC, energy is deposited in the form of scintillation light and ionization electrons. The wavelength spectrum of the scintillation light peaks at 128 nm and therefore needs to be converted into visible light for detection by high-sensitivity cryogenic photosensors (PMTs or SiPMs). The ionization electrons drift toward the anode due to the applied electric field. The X and Y positions are determined at the anode (with a few mm pitch) while the Z positions are determined from the drift time of the ionization electrons after the scintillation light is measured. It is important to note that electronegative impurities such as water and oxygen absorb ionization electrons and therefore high-purity LAr is needed to achieve high sensitivity across the entire detection area.
@@
As shown in the right side of Fig. 1, when a charged antiparticle ($\bar{p},\bar{d},\mathrm{\overline{He}}$) slows down and stops in LAr through ionization energy loss, it forms an exotic atom with the argon nucleus. The exotic atom de-excites, releasing X-rays, and at the end of the transition, the antiparticle annihilates with the nucleus, releasing multiple hadrons, which can also be detected [16]. In summary, the LArTPC can provide 3D-tracking information with deposited energy along the trajectory (dE/dx) for charged particles, allowing for efficient particle detection and the detection of annihilation products (if any) for strong particle identification.
@@ 2 The B23-06 Flight Campaign
Since 2009, science flights that were selected annually by the ISAS/JAXA scientific balloon program advisory committee have been conducted at Taiki Aerospace Research Field (TARF) located in Taiki, Hokkaido, Japan [20]. On September 30, 2022, we submitted a proposal for the GRAMS engineering flight as a JAXA domestic scientific balloon program (PI: Hirokazu Odaka) [21]. In April of 2023, this engineering flight was officially approved as one of the balloon flights for 2023 with two piggyback experiments, designated as B23-06. Flight trajectory for balloons launched from TARF is shown in Figure 3. During the flight season which spans from May to September, the seasonal wind pattern is east wind in the upper stratosphere and strong west-wind jet streams at lower altitudes. This enables a ”boomerang” flight which increases the flight duration and collection of the payload close to the land.
@@
Figure 2: ”Boomerang” flight trajectory from TARF.
@@ 3.2 Detector System
The simulated events of a cosmic ray penetrating through the TPC and a gamma-ray Compton scattering in the middle of the TPC are shown in Figure 10. The outermost channel is labeled ch1, and the innermost channel as ch3. Since there is a grid below the anode, the time at which the signals on each channel start to rise represents the z position of the electron. The y-axis in the plots represents the Integrated Energy (MeV), which was obtained by converting the output of the charge preamplifier from volts to MeV using a conversion factor derived from the expected output of a Minimum Ionizing Particle (MIP) passing through the TPC. For a cosmic ray, the trajectory of the track is shown in the top two plots. For this simulated event, the cosmic ray penetrates through the TPC, hence there will be two signals on ch1 and ch2 and one signal on ch3. For the gamma-ray event, the Compton scattered event would show a point-like event, in which the signal is only present in the channel above the Compton scattering position.
@@ 4 Flight Summary: July 27th 2023
At 8:37 AM on July 26th, LAr was filled into the LAr vessel with enough amount for the flight on the following day. At 2:28 AM on the day of the flight, the final operation check was completed and at 3:55 AM, the balloon was launched. Figure 14 shows the flight trajectory in the left panel and the flight altitude in the right plot with the detector operations for this mission. The GPS data were provided by JAXA/ISAS Balloon Group. As illustrated in the left panel of Figure 14, the balloon follows a boomerang trajectory due to the directions of the jet stream winds at low altitudes being in the east direction and the winds at higher altitudes in the west direction. This boomerang flight enables a longer flight duration and recovery of the gondola closer to the land on the sea. The right plot of Figure 14 shows that the balloon was released at 03:55 AM on July 27th, 2023, after an ascent of 2 hours and 4 minutes, the balloon reached a maximum altitude of 28.9 km, and moved to a level flight at 05:59 AM for 44 minutes. At 06:43 AM, the gondola was released from the balloon and the gondola landed on the sea at 07:07 AM.
@@
The environmental data, including pressure and temperature, along with PMT waveforms, were continuously recorded from the pre-launch phase until all power supplies were shut down. The trigger level for the PMT signal was set to approximately 5 MeV of energy deposition inside the LArTPC fiducial volume which was corresponding to a few cm of charged particle track. To prevent high voltage discharge within the LAr vessel, the high voltage to the cathode of the TPC was turned on 5 minutes after launch and turned off before the release of LAr. The trigger level was lowered for the first 20 minutes of the 40-minute level flight (500 keV for 10 minutes and 100 keV for another 10 minutes) to obtain gamma-ray events with low energy deposits.
@@ 5 LArTPC Data Analysis
As mentioned in section 4, the LArTPC signals were successfully obtained throughout the flight. A total of around 0.5 million events were triggered by the PMT and detected by the LArTPC. The left and right plots in Figure 16 illustrate a cosmic-ray penetration candidate and a gamma-ray Compton scattered candidate obtained during the level flight. The black, red, and blue lines represent the signal on each of the three anode channels. The green line corresponds the common noise and cross-talk component (subtracted from each signal channel). For the cosmic-ray candidate, the signal on ch1 rises at 20 $\mu$s and 60 $\mu$s. Therefore, by considering the drift time of electrons, this could be understood as a cosmic ray entering the TPC from the side 2 cm below the anode and exiting the TPC from another side 6 cm below the anode. For the gamma-ray candidate, the signal on ch3 rises at 20 $\mu$s with no signals on the other channels. Therefore, this could be understood as a gamma-ray Compton scattering that occurred 2 cm below the anode, under the inner pad on the anode.
@@
where $A$ is a scale factor and $R$ is the event rate. The left plot of Fig. 17 shows the distribution of the time difference at an altitude of 5 km (black plots) and the fitting result (red line). The right plot of Figure 17 represents the rates calculated by exponential fitting shown in black. A Geant4-based simulation with input conditions from EXPACS [24] providing angular and energy probability information of each particle at given altitudes was conducted as a comparison to the data. From the Geant4 simulation, the rate of events that deposited more than 5 MeV in the LArTPC is calculated and shown in red. As shown in Figure 17, for both data and the simulation, the event rate increases with the ascent of the balloon up to about 20 km, where the production of secondary particles from primary cosmic rays reaches its peak, known as the shower maximum, before decreasing with further altitude.
Figure 17: (Left) Time distribution between each event at 5 km altitude (black points) and the fitting result (red line). (Right) Event rates at given altitudes are calculated from $\Delta$t distribution (black) Geant4 + EXPACS simulation result (red).
As mentioned in Section 3.1, LAr requires high purity for detecting charge signals on a LArTPC, and the purity needs to be maintained during the flight. In Figure 18 the left and right plots are events obtained at level flight and ground, respectively. As shown in Figure 18, as the particle leaves only a signal on ch3 shown in blue, this event is a penetration event where a particle (most likely a Minimum Ionizing Particle) passes vertically through the TPC. The dotted lines represent how a signal would look with the corresponding LAr purities. It can be seen that the purity was maintained in the flight as the purity was sub-ppb level before and during the flight.
@@ Acknowledgements
We thank the Scientific Ballooning (DAIKIKYU) Research and Operation Group, ISAS, JAXA for the professional support of the eGRAMS flight. We are grateful for the support of Prof. K. Ishimura for the gondola design support. This work was supported by the JSPS Grant-in-Aid for Scientific Research (A) (22H00133), (B) (22H01252), and Challenging Research (Pioneering) (22K18277) in Japan. We also acknowledge support by NASA under Award No.22-APRA22-0128 (80NSSC23K1661) and the Alfred P. Sloan Foundation in the USA.

### v2

@@ 1.2 Cosmic-Ray Antiparticles
Various cosmic particles, including antiparticles such as antiprotons, are incident on the Earth from space. The fluxes of antiprotons have been detected and measured by numerous experiments [18, 13, 16]. The majority of the measured antiproton flux is explained by secondary production in which primary cosmic rays such as protons colliding with interstellar gas: $p_{\mathrm{CR}}+p_{\mathrm{ISM}}\rightarrow\bar{p}+X$. However, a potential excess exists in the low energy region around 1 GeV. While still debated, this excess could be potentially described by the annihilation of dark matter (DM) with a mass around 50 GeV [4]. Despite the high-statistics antiproton measurements, cosmic-ray antideuterons have never been observed. Currently, the only upper limit on the antideuteron flux is placed by BESS Polar-II experiment [15]. When considering the dark matter model that describes the potential antiproton excess, the predicted antideuteron flux at energies below 0.2 GeV/n from dark matter annihilation is significantly larger (S/N $>$ 100) than the expected astrophysical secondary flux. This difference arises due to the kinematic limitations on the secondary production. In contrast, dark matter annihilation lacks these limitations, allowing the production of low-energy antideuterons. Consequently, the search for antideuterons has gained attention as a potential smoking gun for indirect DM search.
@@ 1.3 The GRAMS Experiment
The GRAMS (Gamma-Ray and AntiMatter Survey) experiment is a pioneering next-generation balloon/satellite experiment. It is the first project to utilize a LArTPC for simultaneous astrophysical observations of cosmic MeV gamma-rays and indirect dark matter searches using cosmic antiparticles [1]. Figure 1 illustrates the conceptual design of GRAMS and the detection principle of a LArTPC. As shown in Figure 1, the GRAMS detector consists of a LArTPC surrounded by two layers of Time-of-Flight (ToF) scintillators. For detecting charged particles, the ToFs can provide velocity information of the incoming particle. Meanwhile, vetoing charged particles can provide time windows for detecting gamma rays. When a charged particle enters the LArTPC, energy is deposited in the form of scintillation light and ionization electrons. The wavelength spectrum of the scintillation light peaks at 128 nm and therefore needs to be converted into visible light for detection by high-sensitivity cryogenic photosensors (PMTs or SiPMs). The ionization electrons drift toward the anode due to the applied electric field. The X and Y positions are determined at the anode (with a few mm pitch) while the Z positions are determined from the drift time of the ionization electrons after the scintillation light is measured. It is important to note that electronegative impurities such as water and oxygen absorb ionization electrons and therefore high-purity LAr is needed to achieve high sensitivity across the entire detection area.
@@
As shown in the right side of Fig. 1, when a charged antiparticle ($\bar{p},\bar{d},\mathrm{\overline{He}}$) slows down and stops in LAr through ionization energy loss, it forms an exotic atom with the argon nucleus. The exotic atom de-excites, releasing X-rays, and at the end of the transition, the antiparticle annihilates with the nucleus, releasing multiple hadrons, which can also be detected [2]. In summary, the LArTPC can provide 3D-tracking information with deposited energy along the trajectory (dE/dx) for charged particles, allowing for efficient particle detection and the detection of annihilation products (if any). Combined with velocity information from the ToFs, this enables strong particle identification.
@@ 2 The B23-06 Flight Campaign
Since 2009, science flights that were selected annually by the ISAS/JAXA scientific balloon program advisory committee have been conducted at Taiki Aerospace Research Field (TARF) located in Taiki, Hokkaido, Japan [22]. On September 30, 2022, we submitted a proposal for the GRAMS engineering flight as a JAXA domestic scientific balloon program (PI: Hirokazu Odaka) [11]. In April of 2023, this engineering flight was officially approved as one of the balloon flights for 2023 with two piggyback experiments, designated as B23-06. Flight trajectory for balloons launched from TARF is shown in Figure 3. During the flight season which spans from May to September, the seasonal wind pattern is east wind in the upper stratosphere and strong west-wind jet streams at lower altitudes. This results in a “boomerang” flight, increasing the flight duration and enabling the collection of the payload and the balloon close to the land.
@@
Figure 2: “Boomerang” flight trajectory from TARF.
@@ 3.2 Detector System
The simulated events of a cosmic ray penetrating through the TPC and a gamma-ray Compton scattering in the middle of the TPC are shown in Figure 10. To simulate the output of the charge preamplifiers connected to each anode channel, a step simulation was conducted where a particle was generated and the energy deposited in each step along the track was converted to the output of the charge preamplifier. The outermost channel is labeled ch1, and the innermost channel as ch3. Since there is a grid below the anode, the time at which the signals on each channel start to rise represents the z position of the electron. The y-axis in the plots represents the Integrated Energy (MeV), which was obtained by converting the output of the charge preamplifier from volts to MeV using a conversion factor derived from the expected output of a Minimum Ionizing Particle (MIP) passing through the TPC. For a cosmic ray, the trajectory of the track is shown in the top two plots. For this simulated event, the cosmic ray penetrates through the TPC, hence there will be two signals on ch1 and ch2 and one signal on ch3. For the gamma-ray event, the Compton scattered event would show a point-like event, in which the signal is only present in the channel above the Compton scattering position.
@@ 4 Flight Summary: July 27th 2023
At 8:37 AM on July 26th, LAr was filled into the LAr vessel with enough amount for the flight on the following day. Until 12:00 AM, the LArTPC was operated to obtain cosmic muon events at the ground for calibration measures. At 2:28 AM on the day of the flight, the final operation check was completed and at 3:55 AM, the balloon was launched. Figure 14 shows the flight trajectory in the left panel and the flight altitude in the right plot with the detector operations for this mission. The GPS data were provided by JAXA/ISAS Balloon Group. As illustrated in the left panel of Figure 14, the balloon follows a boomerang trajectory due to the directions of the jet stream winds at low altitudes being in the east direction and the winds at higher altitudes in the west direction. This boomerang flight enables a longer flight duration and recovery of the gondola closer to the land on the sea. The right plot of Figure 14 shows that the balloon was released at 03:55 AM on July 27th, 2023, after an ascent of 2 hours and 4 minutes, the balloon reached a maximum altitude of 28.9 km, and moved to a level flight at 05:59 AM for 44 minutes. At 06:43 AM, the gondola was released from the balloon and the gondola landed on the sea at 07:07 AM.
@@
The environmental data, including pressure and temperature, along with PMT waveforms, were continuously recorded from the pre-launch phase until all power supplies were shut down. The trigger level for the PMT signal was set to approximately 5 MeV of energy deposition inside the LArTPC active volume which was corresponding to a few cm of charged particle track. To prevent high voltage discharge within the LAr vessel, the high voltage to the cathode of the TPC was turned on 5 minutes after launch and turned off before the release of LAr. The trigger level was lowered for the first 20 minutes of the 40-minute level flight (500 keV for 10 minutes and 100 keV for another 10 minutes) to obtain gamma-ray events with low energy deposits.
@@ 5 LArTPC Data Analysis
As mentioned in section 4, the LArTPC signals were successfully obtained throughout the flight. A total of around 0.5 million events were triggered by the PMT and detected by the LArTPC. The left and right plots in Figure 16 illustrate a cosmic-ray penetration candidate and a gamma-ray Compton scattered candidate obtained during the level flight. The black, red, and blue lines represent the signal on each of the three anode channels. The green line corresponds to the common noise component (subtracted from each signal channel). For the cosmic-ray candidate, the signal on ch1 rises at 20 $\mu$s and 60 $\mu$s. Therefore, by considering the drift time of electrons, this could be understood as a cosmic ray entering the TPC from the side 2 cm below the anode and exiting the TPC from another side 6 cm below the anode. For the gamma-ray candidate, the signal on ch3 rises at 20 $\mu$s with no signals on the other channels. Therefore, this could be understood as a gamma-ray Compton scattering that occurred 2 cm below the anode, under the inner pad on the anode.
@@
where $A$ is a scale factor and $R$ is the event rate. The left plot of Fig. 17 shows the distribution of the time difference at an altitude of 5 km (blue dots), 20 km (black dots), and the corresponding fitting results (red lines). The right plot of Figure 17 represents the rates calculated by exponential fitting shown in black. A Geant4-based simulation with input conditions from EXPACS [24] providing angular and energy probability information of each particle at given altitudes was conducted as a comparison to the data. From the Geant4 simulation, the rate of events that deposited more than 5 MeV in the LArTPC is calculated and shown in red. As shown in Figure 17, for both data and the simulation, the event rate increases with the ascent of the balloon up to about 20 km, where the production of secondary particles from primary cosmic rays reaches its peak, known as the shower maximum, before decreasing with further altitude.
Figure 17: (Left) Time distribution between each event at 5 km altitude (blue dots), 20 km altitude (black dots), and the fitting results (red lines). (Right) Event rates at given altitudes are calculated from $\Delta$t distribution (black) Geant4 + EXPACS simulation result (red).
As mentioned in Section 3.1, LAr requires high purity for detecting charge signals on a LArTPC, and the purity needs to be maintained during the flight. In Figure 18 the left and right plots are events obtained at level flight and ground, respectively. As shown in Figure 18, as the particle leaves only a signal on ch3 shown in blue, this event is a penetration event where a particle (most likely a Minimum Ionizing Particle) passes vertically through the TPC. The dotted lines represents fitted signal functions with the corresponding LAr impurities when the integrated energy of the signal is kept as a constant parameter. It can be seen that the purity was maintained in the flight as the purity was sub-ppb level before and during the flight.
@@ Acknowledgements
We thank the Scientific Ballooning (DAIKIKYU) Research and Operation Group, ISAS, JAXA for the professional support of the eGRAMS flight. We are grateful for the support of Prof. K. Ishimura for the gondola design support. Also, we are grateful for the support of Prof. Y. Makida for providing the absolute pressure valve and for consultations on the safety design. This work was supported by the JSPS Grant-in-Aid for Scientific Research (A) (22H00133), (B) (22H01252), and Challenging Research (Pioneering) (22K18277) in Japan. We also acknowledge support by NASA under Award No.22-APRA22-0128 (80NSSC23K1661) and the Alfred P. Sloan Foundation in the USA.